Jody's older brother is only three years less than twice jody's age. If sum of thier age is 30, then what is jody's brother age

Answer by addingup(3677)

brother: b
Jody: j
brother is three years less than twice jody's age:
b = 2j-3
--------------
j+b = 30 substitute for brother:
j+2j-3 = 30
3j = 33
j = 11
and
b = 2j-3
b = 2(11)-3
b = 19